# fMRI-3D: A Comprehensive Dataset for Enhancing fMRI-based 3D Reconstruction

### v1

@@ Abstract
Reconstructing 3D visuals from functional Magnetic Resonance Imaging (fMRI) data, introduced as Recon3DMind in our conference work [1], is of significant interest to both cognitive neuroscience and computer vision. To advance this task, we present the fMRI-3D dataset, which includes data from 15 participants and showcases a total of 4,768 3D objects. The dataset comprises two components: fMRI-Shape, previously introduced and accessible at https://huggingface.co/datasets/Fudan-fMRI/fMRI-Shape, and fMRI-Objaverse, proposed in this paper and available at https://huggingface.co/datasets/Fudan-fMRI/fMRI-Objaverse. fMRI-Objaverse includes data from 5 subjects, 4 of whom are also part of the Core set in fMRI-Shape, with each subject viewing 3,142 3D objects across 117 categories, all accompanied by text captions. This significantly enhances the diversity and potential applications of the dataset. Additionally, we propose MinD-3D, a novel framework designed to decode 3D visual information from fMRI signals. The framework first extracts and aggregates features from fMRI data using a neuro-fusion encoder, then employs a feature-bridge diffusion model to generate visual features, and finally reconstructs the 3D object using a generative transformer decoder. We establish new benchmarks by designing metrics at both semantic and structural levels to evaluate model performance. Furthermore, we assess our model’s effectiveness in an Out-of-Distribution setting and analyze the attribution of the extracted features and the visual ROIs in fMRI signals. Our experiments demonstrate that MinD-3D not only reconstructs 3D objects with high semantic and spatial accuracy but also deepens our understanding of how human brain processes 3D visual information. Project page at: https://jianxgao.github.io/MinD-3D.
@@
FMRI decoding, 3D vision, Dataset, Diffusion model.1 Introduction
@@
Existing methods primarily focus on reconstructing 2D visual information, but the human visual system extends far beyond merely processing flat images. It possesses the extraordinary ability to transform 2D projections into rich 3D representations. This complex mechanism allows us to perceive the world in depth, recognizing attributes like size, distance, and spatial depth. In contrast to previous studies, our research centers on modeling the brain’s 3D visual capabilities. We introduce a new task, called Recon3DMind (Reconstructing 3D Objects from Mind), which leverages advanced computer vision techniques to decode and reconstruct the 3D visual information perceived by the brain from fMRI signals. This task goes beyond merely extracting semantic features, incorporating spatial and structural dimensions that are essential for a comprehensive understanding of 3D vision.
@@
Leveraging our carefully curated fMRI-3D dataset, we introduce an innovative and efficient three-stage framework called MinD-3D to extract both spatial structure and semantic features from multi-frame fMRI signals and reconstruct the corresponding 3D visual stimuli.
@@
In the first stage, we utilize an encoder [6] pre-trained on NSD [16] dataset to extract spatial features from the fMRI data. These features are then aggregated across multiple frames using a feature aggregation module. To ensure the biological relevance and effectiveness of these extracted features, we align them with the visual-spatial characteristics of the corresponding objects using contrastive learning loss. This alignment ensures that the extracted features are both accurate and biologically meaningful in relation to the brain’s visual processing mechanisms.
The second stage focuses on decoding the brain’s visual activity from the fMRI data. We achieve this through a transformer-based diffusion model trained in feature space, conditioned on the fMRI-derived features. This conditioning enables us to progressively generate accurate visual representations, translating complex brain activity into understandable visual data.
In the final stage, we focus on reconstructing the 3D models as they are perceived by the human brain. To do this, we train a latent-adapted Argus [17] model, which uses the visual features generated in the second stage along with the initial fMRI features to create 3D objects. This process mimics how the brain’s visual cortex operates, providing a simulated insight into the brain’s perception and processing of 3D objects.
@@
To measure the effectiveness of our model, we design new benchmarks that evaluate its performance at both the semantic and structural levels. These metrics comprehensively assess the model’s ability to generate 3D representations that are structurally and semantically accurate. We evaluate our model under both standard and Out-of-Distribution settings, and it consistently outperforms baseline models. Additionally, we conduct in-depth analyses of our fMRI-3D dataset and the features extracted by MinD-3D, exploring how the brain perceives different angles, objects, and semantic information within specific ROIs (regions of interest). We further validate the relevance of our model’s features by correlating them with brain regions, confirming that the representations it produces are consistent with the brain’s visual information processing.
@@
This paper expands upon our preliminary conference work, and we summarize the key contributions as follows:
@@
We introduce fMRI-Objaverse, a large-scale extension of the original dataset. Together with fMRI-Shape, we collectively refer to these datasets as fMRI-3D, designed to support various experimental setups and advance research in the field.
@@
We establish new benchmarks for the task of 3D visual reconstruction from human brain data.
@@
We conduct extensive experiments to analyze the contributions of our proposed dataset in decoding fMRI signals, further validating the effectiveness of MinD-3D.
@@ 2.2 Diffusion Models
3 Experimental Designs and Curated Dataset
In this section, we detail the procedures for collecting the proposed fMRI-3D dataset, which consists of two components: fMRI-Shape and fMRI-Objaverse. The scale of fMRI-3D is compared to other benchmark datasets, including NSD [16], BOLD5000 [35], GOD [18], and Video-fMRI [19], as illustrated in Fig. 4. Specific details about fMRI-Shape and fMRI-Objaverse are provided in Tab. I. For all experiments, written informed consent was obtained from each participant, and the study was approved by the ethical review board. To better illustrate brain activation patterns and demonstrate the utility of the fMRI-3D dataset, we analyze and visualize responses to three distinct objects across six subjects, as shown in Fig. 3. Note that only voxels with activation levels above the 50th percentile are displayed. We also compute the variation across subjects and objects, with red and blue regions indicating higher activation values, respectively, reflecting areas in the human brain sensitive to the stimuli. This visualization highlights significant individual differences in brain activation across subjects, which are more pronounced than the variations in responses to different objects. These findings emphasize the inherent challenges and underscore the importance of the AP and APAC settings. All participants had normal or corrected-to-normal vision. The fMRI-3D dataset will be made publicly available to support further research in Recon3DMind.
| Participant | Males/Females | Category | Objects | Frames | |
@@ 3 Experimental Designs and Curated Dataset
3.1 fMRI-Shape
As illustrated in the middle part of Fig. 6, individual differences among participants pose significant challenges for generalization. The AP and APAC sets are crucial for OOD testing and will serve as important benchmarks for assessing the generalization capability of 3D decoding models.
@@
FMRI-Objaverse, as partially shown in Fig. 5, includes data from five participants, all of whom were unfamiliar with the study. Four participants (2 males and 2 females, aged 22 to 26, identified as Nos. 1, 6, 7, and 8) overlap with the core fMRI-Shape dataset, providing an important extension in terms of diversity and scale. Additionally, the fifth participant (No. 15), a 22-year-old female, was included to further expand the dataset. To enhance our dataset, we selected 3,142 objects from the top 117 object categories in Objaverse [15], based on a subset filtered by LGM [27] and enriched with text descriptions from Cap3D [40]. Unlike in fMRI-Shape, each 3D object in this dataset was rendered into 384 frames, generating a 6.4-second video at 48 fps using Blender. Each participant spent approximately 8 hours in experimental sessions, divided into 53 sessions. During each session, participants viewed 60 videos in a randomized order, except for the last session, with 1.6-second rest intervals between each pair of objects. To prevent low-quality data due to visual fatigue, we randomly reversed the rotation direction for 40% of the selected objects. All objects were presented once to each participant. (Note: Our MRI machine samples data every 800ms, so we selected 6.4-second videos with a 1.6-second rest period between them.) This extension supports further multimodal experiments and applications. The objects in Objaverse contain more detail and a wider variety of categories, and the higher fps videos present stronger visual effects, posing a significant challenge for reconstructing them in fMRI-Objaverse. The dataset is available for download at: https://huggingface.co/datasets/Fudan-fMRI/fMRI-Objaverse.
@@ 3.3 Data Acquisition and Preprocessing
T1 and fMRI data were acquired in a 3T scanner and a 32-channel RF head coil. T1-weighted data were scanned using MPRAGE sequence (0.8-mm isotropic resolution, TR=2500ms, TE=2.22ms, flip angle ). Functional data were scanned using gradient-echo EPI at 2-mm isotropic resolution with whole-brain coverage (TR=800ms, TE=37ms, flip angle , multi-band acceleration factor 8). The sampling frequency of the 3T scanner is 1.25Hz, so each video segment corresponds to a total of 10 frames of task-state fMRI signals.
@@
Preprocessing was performed using fMRIPrep [41, 42]. Following [5], the preprocessed functional data in 32k_fs_LR surface space were converted into 2D images and utilized for further analysis. Given the delay of the BOLD signal by 6 seconds, we applied z-scoring to the data points across every vertex within each run, incorporating a 6.4-second lag. These normalized values were then projected onto 1023 × 2514 pixel 2D images using pycortex. For analysis, Regions of Interest (ROIs) were selected from the Human Connectome Project Multi-Modal Parcellation (HCP-MMP) atlas in the 32k_fs_LR space. These ROIs included areas such as “V1, V2, V3, V3A, V3B, V3CD, V4, LO1, LO2, LO3, PIT, V4t, V6, V6A, V7, V8, PH, FFC, IP0, MT, MST, FST, VVC, VMV1, VMV2, VMV3, PHA1, PHA2, PHA3”. Subsequently, the ROIs were converted into a 256 × 256 image, as illustrated in the right part of Fig. 6.
@@ 4 Problem Setup
We begin with the acquisition of a multi-frame fMRI signal, denoted as , where (with or ). These signals correspond to both a 3D object mesh, , and a video, , with (where or frames), which the subject observes. The task requires an efficient encoder, , capable of extracting both spatial structural and semantic features from the fMRI signal. It is important to note that while a single frame of fMRI data is sufficient to extract semantic information for 2D image reconstruction, reconstructing 3D structures requires additional spatial structural features. Therefore, multiple frames of fMRI data are input to capture these comprehensive spatial features from the spatio-temporal signals. Mathematically, this is expressed as: .
After feature extraction, a powerful decoder is used to reconstruct the original 3D mesh, , based on the extracted feature : . Thus, our model can be succinctly described as , where the transformation is represented as .
@@ 5 Method
5.1 Overview
As depicted in Fig. 7, our MinD-3D model comprises three main components: 1) Neuro-Fusion Encoder: This part is responsible for extracting both semantic and structural features from the fMRI frames. 2) Feature Bridge Diffusion Model: This module generates visual features based on the features extracted by the Encoder. 3) Latent Adapted Decoder: Finally, this component recovers 3D objects, using both the fMRI features and the generated visual features.
5.2 Neuro-Fusion Encoder
Drawing from the potent Masked Image Model (MIM) [43, 44, 6], we adopt an auto-encoder structure to extract meaningful fMRI features, trained on the UKB dataset [45]. Our encoder, termed the Neuro-Fusion Encoder (NFE), employs the encoder of LEA [6] as , along with a Feature Aggregation module (), to transform fMRI frames into embeddings. Inspired by contrastive learning [46, 47], we then align these embeddings with the vision space through CLIP [46].
@@ 5.2 Neuro-Fusion Encoder
Initially, we process each fMRI frame in parallel to obtain the spatial fMRI embeddings:
Then, we use the Feature Aggregation module to aggregate the fMRI embeddings into fMRI latent feature :
To align the fMRI latent feature with the vision space, we employ the ViT-B/32 CLIP vision encoder and here to encode video frames into vision embeddings. During training, we randomly select n frames as and compute the average vision embeddings as the visual latent feature :
where n is the number of selected video frames. Then we utilize the CLIP loss to ensure the alignment between fMRI features and visual features :
During training, we optimize the neuro-fusion encoder initialized with pre-trained weights, while the CLIP vision encoder remains frozen. It’s worth noting that the CLIP vision encoder and images are just used for training, we will drop it during inference. By adopting this approach, we will get the fMRI features and embedding aligned with the visual space. The fMRI features will be used as the conditional information for the feature bridge diffusion model.
5.3 Feature Bridge Diffusion Model
This section describes the Neural-Visual Synthesis process via a transformer-based diffusion model, herein referred to as the Feature Bridge Diffusion Model (FBDM), inspired by the hierarchical design proposed in [48]. The FBDM’s primary role is to bridge fMRI latent features with their visual counterparts , effectively transforming neurological signals into visual representations. In the forward diffusion process, we treat the visual latent feature as and introduce Gaussian noise into it across 100 timesteps. This approach modifies the standard 1000-timestep model to align with the settings in Mind-eye [4]:
During the reverse diffusion process, FBDM is tasked with predicting and reversing the introduced noise to reconstruct the latent visual feature. This is achieved by estimating the noise at each timestep, conditioned on the fMRI latent feature :
Differing from U-Net based models, our approach utilizes a transformer to leverage the self-attention mechanism for integrating the conditional information . This is done by embedding the timestep , the noise-perturbed feature , and the conditional fMRI feature into a unified representation, further augmented with a learnable token to capture the target vision features accurately. The objective of FBDM training is to minimize the discrepancy between the actual noise and its prediction:
Upon training completion, the model discards the initial visual latent feature and enters a generative phase. It begins with pure noise and progressively refines this through the reverse diffusion process to generate the predicted visual latent feature . This process effectively translates the aligned fMRI latent features into visual representations, serving as essential conditional information for the subsequent latent-adapted decoding phase.
5.4 Latent Adapted Decoder
The Latent Adapted Decoder (LAD) leverages Argus [17], a state-of-the-art 3D mesh generation method, as its backbone. This component is tasked with generating the Vector Quantized (VQ) latent representation of target 3D objects from conditional information processed through a generative auto-regressive transformer. The resulting 3D objects are then reconstructed using the VQ decoder.
To tailor Argus for processing fMRI input and to retain its generative efficiency, we employ adapter tuning during training. This involves integrating the adapter layers within the transformer block. Following classic methods [49, 50], we apply two feed-forward linear layers to effectively adapt the self-attention outputs. The adapter layer is systematically inserted at every transformer blocks, ensuring that the model remains adaptable and efficient. The decoder formulates the VQ latent vector for the target object as follows:
Subsequently, the 3D VQ-decoder decodes this latent vector into the final 3D mesh representation :
The decoder aims to minimize the negative log-likelihood loss during optimization:
For training, the decoder is fine-tuned from pre-trained weights, with the trained neuro-fusion encoder and the feature bridge diffusion model. This strategy preserves Argus’s generative capabilities while adapting to new inputs and minimizing the computational overhead associated with training.
| Methods | Semantic-Level | Structure-Level | |||||
| 2-way | 10-way | LPIPS | SSIM | FPD | CD | EMD | |
| LEA-3D [6] | 0.787 | 0.371 | 0.527 | 0.562 | 4.229 | 2.291 | 5.347 |
| fMRI-PTE-3D [5] | 0.815 | 0.392 | 0.433 | 0.694 | 3.571 | 1.992 | 4.621 |
| w/o Both | 0.789 | 0.367 | 0.479 | 0.616 | 3.694 | 2.205 | 5.073 |
| w/o Diffusion | 0.801 | 0.385 | 0.423 | 0.669 | 3.526 | 2.071 | 4.625 |
| w/o Contrastive | 0.823 | 0.419 | 0.319 | 0.701 | 3.315 | 1.826 | 4.027 |
| MinD-3D (full) | 0.839 | 0.432 | 0.230 | 0.734 | 3.157 | 1.742 | 3.833 |

### v3

# MinD-3D++: Advancing fMRI-Based 3D Reconstruction with High-Quality Textured Mesh Generation and a Comprehensive Dataset

@@ Abstract
Reconstructing 3D visuals from functional Magnetic Resonance Imaging (fMRI) data, introduced as Recon3DMind, is of significant interest to both cognitive neuroscience and computer vision. To advance this task, we present the fMRI-3D dataset, which includes data from 15 participants and showcases a total of 4,768 3D objects. The dataset consists of two components: fMRI-Shape, previously introduced and available at https://huggingface.co/datasets/Fudan-fMRI/fMRI-Shape, and fMRI-Objaverse, proposed in this paper and available at https://huggingface.co/datasets/Fudan-fMRI/fMRI-Objaverse. fMRI-Objaverse includes data from 5 subjects, 4 of whom are also part of the core set in fMRI-Shape. Each subject views 3,142 3D objects across 117 categories, all accompanied by text captions. This significantly enhances the diversity and potential applications of the dataset. Moreover, we propose MinD-3D++, a novel framework for decoding textured 3D visual information from fMRI signals. The framework evaluates the feasibility of not only reconstructing 3D objects from the human mind but also generating, for the first time, 3D textured meshes with detailed textures from fMRI data. We establish new benchmarks by designing metrics at the semantic, structural, and textured levels to evaluate model performance. Furthermore, we assess the model’s effectiveness in out-of-distribution settings and analyze the attribution of the proposed 3D pari fMRI dataset in visual regions of interest (ROIs) in fMRI signals. Our experiments demonstrate that MinD-3D++ not only reconstructs 3D objects with high semantic and spatial accuracy but also provides deeper insights into how the human brain processes 3D visual information. Project page: https://jianxgao.github.io/MinD-3D.
@@
FMRI decoding, 3D vision, Dataset, Diffusion model.
1 Introduction
@@
Existing methods primarily focus on reconstructing 2D visual information, but the human visual system extends far beyond merely processing flat images. The human brain appears capable of interpreting 2D projections in ways that support the perception of 3D structure [8, 9]. This complex mechanism allows us to perceive the world in depth, recognizing attributes like size, distance, and spatial depth. In contrast to previous studies, our research centers on modeling the brain’s 3D visual capabilities. We introduce a new task, called Recon3DMind (Reconstructing 3D Objects from Mind), which leverages advanced computer vision techniques to decode and reconstruct the 3D visual information perceived by the brain from fMRI signals. This task goes beyond merely extracting semantic features, incorporating spatial and structural dimensions that are essential for a comprehensive understanding of 3D vision.
@@
Leveraging our carefully curated fMRI-3D dataset, we introduce MinD-3D++, a novel framework which generates textured 3D visual stimuli directly from fMRI signals for the first time. The framework comprises two main steps: (1) extracting features from multi-frame fMRI signals, and (2) generating corresponding multi-view images and subsequently synthesizing 3D objects.
@@
During feature extraction, we first use a transformer-based encoder [5] pre-trained on the NSD dataset [18] to extract spatial features from the fMRI data. These features are then aggregated across multiple frames through a feature aggregation module. To maintain biological relevance and ensure the effectiveness of the extracted features, we align them with both the visual and textual representations of the corresponding objects. This alignment is achieved by applying a contrastive learning loss on the class token within the encoder’s blocks.
Once the features are aligned, we leverage the strong generative capabilities of diffusion models to accurately capture object appearance. Specifically, we adopt a multi-view diffusion model in which the extracted features and class token serve as conditional inputs to a pre-trained diffusion model. Through LoRA-based fine-tuning of the attention layers, we enable the model to incorporate fMRI-derived features. This process yields six-view images that faithfully reflect the original 3D stimuli. Finally, using a pre-trained model, we synthesize a textured 3D mesh from these multi-view images, completing the end-to-end generation of 3D visual stimuli.
@@
To evaluate the effectiveness of our model, we design new benchmarks that measure performance across semantic, structural, and textured levels. These metrics provide a comprehensive assessment of the model’s ability to generate 3D representations that are both structurally and semantically accurate. We test our model in both standard and out-of-distribution settings, where it consistently outperforms baseline models. Additionally, we conduct in-depth analyses of the fMRI-3D dataset and the features extracted by MinD-3D. These analyses explore how the brain perceives different angles, objects, and semantic information within specific regions of interest (ROIs). We further validate the biological relevance of our model’s features by correlating them with brain regions, demonstrating that the generated representations align with the brain’s visual information processing.
@@
This paper builds upon our preliminary conference work, and we summarize the key contributions as follows:
@@
We introduce fMRI-Objaverse, a large-scale extension of the original dataset. Together with fMRI-Shape, we collectively refer to these datasets as fMRI-3D, which is designed to support various experimental setups and advance research in the field.
We propose an improved framework, MinD-3D++, for the first time, capable of reconstructing textured meshes from fMRI signals, representing a significant advancement in decoding 3D representations from human mind.
@@
We establish a comprehensive benchmark for the task of 3D visual reconstruction from human brain data.
@@
We conduct extensive experiments to analyze the contributions of our proposed dataset to decoding fMRI signals, further validating the effectiveness of MinD-3D++.
@@ 2.2 Diffusion Models
3 Curated Dataset
In this section, we detail the procedures for collecting the proposed fMRI-3D dataset, which consists of two components: fMRI-Shape and fMRI-Objaverse. The scale of fMRI-3D is compared to other benchmark datasets, including NSD [18], BOLD5000 [37], GOD [19], and Video-fMRI [20], as illustrated in Fig. 4. Specific details about fMRI-Shape and fMRI-Objaverse are provided in Tab. I. For all experiments, written informed consent was obtained from each participant, and the study was approved by the ethical review board. To better illustrate brain activation patterns and demonstrate the utility of the fMRI-3D dataset, we analyze and visualize responses to three distinct objects across six subjects, as shown in Fig. 3. Note that only voxels with activation levels above the 50th percentile are displayed. We also compute the variation across subjects and objects, with red and blue regions indicating higher activation values, respectively, reflecting areas in the human brain sensitive to the stimuli. This visualization highlights significant individual differences in brain activation across subjects, which are more pronounced than the variations in responses to different objects. These findings emphasize the inherent challenges and underscore the importance of the Across-Person (AP) and Across-Person & Across-Class (APAC) settings. All participants 111Acknowledgement: We thank all volunteers for participating in our fMRI experiments. We are especially grateful to Yansong Bai, Ruiying Chen, Yachen Gao, Junwei Liu, Yunheng Li, Tiran Li, Tongying Pan, Yao Xiao, and Yuxi Zheng for their additional support during the data collection process. had normal or corrected-to-normal vision. The fMRI-3D dataset will be made publicly available to support further research in Recon3DMind.
3.1 fMRI-Shape
As illustrated in Fig. 3, and the middle part of Fig. 7, individual differences among participants pose significant challenges for generalization. The AP and APAC sets are crucial for OOD testing and will serve as important benchmarks for assessing the generalization capability of 3D decoding models.
P. | Male/Female | Category | Obj | Frames | |
@@ 3 Curated Dataset
FMRI-Objaverse, as partially shown in Fig. 6, includes data from five participants, all of whom were unfamiliar with the study. Four participants (2 males and 2 females, aged 22 to 26, identified as Nos. 1, 6, 7, and 8) overlap with the core fMRI-Shape dataset, providing an important extension in terms of diversity and scale. Additionally, the fifth participant (No. 15), a 22-year-old female, was included to further expand the dataset. To enhance our dataset, we selected 3,142 objects from the top 117 object categories in Objaverse [17], based on a subset filtered by LGM [29] and enriched with text descriptions from Cap3D [42]. Unlike in fMRI-Shape, each 3D object in this dataset was rendered into 384 frames, generating a 6.4-second video at 48 fps using Blender. Each participant spent approximately 8 hours in experimental sessions, divided into 53 sessions. During each session, participants viewed 60 videos in a randomized order, except for the last session, with 1.6-second rest intervals between each pair of objects. To prevent low-quality data due to visual fatigue, we randomly reversed the rotation direction for 40% of the selected objects. All objects were presented once to each participant. (Note: Our MRI machine samples data every 800ms, so we selected 6.4-second videos with a 1.6-second rest period between them.) This extension supports further multimodal experiments and applications. The objects in Objaverse contain more details and a wider variety of categories, and the higher fps videos present stronger visual effects, posing a significant challenge for reconstructing them in fMRI-Objaverse. The dataset is available for download at: https://huggingface.co/datasets/Fudan-fMRI/fMRI-Objaverse.
@@ 3.3 Data Acquisition and Preprocessing
The T1 and fMRI data were acquired in a 3T scanner and a 32-channel RF head coil. T1-weighted data were scanned using MPRAGE sequence (0.8-mm isotropic resolution, TR=2500ms, TE=2.22ms, flip angle ). Functional data were scanned using gradient-echo EPI at 2-mm isotropic resolution with whole-brain coverage (TR=800ms, TE=37ms, flip angle , multi-band acceleration factor 8). The sampling frequency of the 3T scanner is 1.25Hz, so each video segment corresponds to a total of 10 frames of task-state fMRI signals.
@@
Preprocessing was performed using fMRIPrep [43, 44]. Following [4], the preprocessed functional data in 32k_fs_LR surface space were converted into 2D images and utilized for further analysis. Given the delay of the BOLD signal by 6 seconds, we applied z-scoring to the data points across every vertex within each run, incorporating a 6.4-second lag. These normalized values were then projected onto 1023 × 2514 pixel 2D images using pycortex. For analysis, Regions of Interest (ROIs) were selected from the Human Connectome Project Multi-Modal Parcellation (HCP-MMP) atlas in the 32k_fs_LR space. These ROIs included areas such as “V1, V2, V3, V3A, V3B, V3CD, V4, LO1, LO2, LO3, PIT, V4t, V6, V6A, V7, V8, PH, FFC, IP0, MT, MST, FST, VVC, VMV1, VMV2, VMV3, PHA1, PHA2, PHA3, TE2p”. Subsequently, the ROIs were converted into a 256 × 256 image, as illustrated in the right part of Fig. 7.
@@ 4 Problem Setup
We begin with the acquisition of a multi-frame fMRI signal, denoted as , where (with or ). These signals correspond to both a 3D object mesh , and a video , with (where or frames), which the subject observes. The task requires an efficient encoder, , capable of extracting both spatial structural and semantic features from the fMRI signal. It is important to note that while a single frame of fMRI data is sufficient to extract semantic information for 2D image reconstruction, reconstructing 3D structures requires additional spatial structural features. Therefore, multiple frames of fMRI data are input to capture these comprehensive spatial features from the spatio-temporal signals.
After feature extraction, a powerful decoder is used to reconstruct the original 3D mesh , based on the extracted feature : . Thus, our model can be succinctly described as , where the transformation is represented as .
@@ 5 Method
5.1 Preliminary
Our conference model, MinD-3D [15], shown in Fig. 9, demonstrates the feasibility of reconstructing 3D meshes from human brain data. MinD-3D combines a neuro-fusion encoder for extracting features from fMRI frames, a feature-bridged diffusion model for generating visual features from these fMRI signals, and a latent-adapted decoder based on the Argus 3D shape generator for reconstructing 3D objects. This integrated system effectively aligns and translates brain signals into accurate 3D visual representations.
Despite the success of MinD-3D, one important aspect remains unexplored: the reconstruction of 3D objects with texture. Inspired by the strong appearance-generation capabilities of 2D diffusion models, we propose a novel approach for exploring textural reconstruction. This approach involves enhancing the encoder and adopting a more robust, multi-view-based 3D generation model. Our goal is to use the diffusion model to generate both the appearance and texture of 3D objects. This represents the first attempt to generate textured meshes from fMRI data and expands the range of object categories that can be decoded from fMRI into 3D representations. Therefore, we introduce the improved framework, MinD-3D++, as illustrated in Fig. 8.
5.2 Neuro Align Encoder
We used the same encoder architecture as in MinD-3D. In the original work LEA [5], the encoder utilized a class token that was trained to reconstruct the complete raw fMRI signal by masking all spatial tokens. This class token is informative and powerful, and in this improved model, we focus on utilizing the class token and apply contrastive learning to it, using it as part of the conditional information for the subsequent diffusion model.
@@ 5.2 Neuro Align Encoder
We still process each fMRI frame in parallel to obtain both the spatial fMRI embeddings and the class tokens:
where .
For spatial information from the fMRI signals, we still employ an aggregation module to obtain the latent fMRI feature:
To enhance the information from the class token , we align it with visual and additional textual feature spaces. Specifically, since we use frames of fMRI as input, we average the class tokens over all frames:
Next, to improve the performance of contrastive learning, we use ViT-H CLIP to extract multi-view image and text features. For the multi-view images, we randomly select one of the six rendered views and compute the visual feature using the CLIP vision encoder , For text, we directly use the CLIP text encoder to extract the feature:
We then calculate the contrastive learning losses between the fMRI features extracted from half of the encoder’s transformer blocks and the visual and textual features, in order to enhance the quality of the extracted features.
Thus the constrastive loss for the encoder is defined as:
During training, we optimize the neuro-align encoder (initialized with pre-trained weights), while keeping the CLIP encoders frozen. It is important to note that the CLIP encoders and , along with the images and text , are used only during training and are discarded during inference.
This approach enables us to align the fMRI features and embeddings with both the visual and textual spaces. These aligned fMRI features will then serve as the conditional information for the multi-view diffusion model.
5.3 3D Generation
To fully utilize the fMRI signal features and generate textured 3D objects, we design a pipeline that generates multi-view images and synthesizes them into 3D models. Let represent the fMRI features and the embeddings derived from the neuro align encoder. The objective is to generate multi-view images conditioned on these fMRI signals.
In our model, and serve as conditional latent inputs for the cross-attention mechanism in the Multi-View Diffusion Model. To improve training efficiency and fully utilize the generation capabilities of the pretrained model, we employ a Low-Rank Adaptation (LoRA) [45] fine-tuning strategy. Specifically, LoRA is applied to the projection layers of the query and value matrices in both the cross-attention and self-attention modules, while keeping the rest of the model parameters frozen.
To train the model, we prepare GT multi-view images by rendering 3D objects into six target images at a resolution of with a white background in blender. The poses of the six images are defined by interleaving absolute elevations of and , combined with azimuths relative to the query image, starting at and increasing by for each subsequent pose. The pretrained diffusion model [46] generates images, which represent six multi-view images arranged in a grid. Each of these images is resized to for processing.
During the reverse diffusion process, the Multi-View Diffusion Model estimates the noise at each timestep , conditioned on the fMRI feature and embeddings :
The training objective minimizes the discrepancy between the predicted noise and the true noise through the following loss function:
Then, the loss function for our model is :
After obtaining the multi-view images of the 3D object, we employ the off-the-shelf sparse-view LRM [46] method to generate the final 3D textured mesh.